# Learning to Generate Unit test via Adversarial Reinforcement Learning

### v2

@@ A.1 Grouped Relative Policy Optimization
A.2 Additional Experiment
Ablation on validity reward
@@ A.2 Additional Experiment
In order to demonstrate the effectiveness of the validity reward and its design choice (see Equation 2), we conduct an ablation study comparing the following variants of the UTRL, (1) UTRL trained without validity reward (i.e., w/o $R_{\text{valid}}$), and (2) UTRL with validity reward but without clipped normalization in validity reward term, where the validity reward is defined as a ratio of valid test cases among the generated test cases (i.e., w/o clipping). Figure 10 shows the number of test cases and the ratio of valid test cases in the generated unit test over training steps. UTRL without $R_{\text{valid}}$ results in the generation of unit tests with more than 50% of invalid test cases, while UTRL results in unit tests containing less than 35% of invalid test cases. Furthermore, in the case of UTRL without clipping, we observe that the model collapses to generate an extremely small number of trivial test cases in order to maximize the ratio of the valid test cases, hindering the learning progress toward generating comprehensive unit tests. These results demonstrates the effectiveness of the design choice of UTRL.
Best-of-N improvement measured with additional code generators
@@
Qwen3-4B + UTRL 0.536 +0.207 14.2 +7.0 0.551 +0.353 16.7 +10.2
Qwen3-14B + UTRL 0.548 +0.219 15.0 +7.8 0.557 +0.359 17.5 +11.0
GT (upperbound) 0.650 +0.304 21.0 +13.1 0.680 +0.278 24.4 +13.7
Table 3: Best-of-N improvement measured using Qwen3-4B and GPT-4o as code generator LLM. UTRL shows the highest best-of-N improvement compared to baselines. Here, UTRL denotes the variant of UTRL using 4.5K programming tasks for training, for fair comparison to CURE.
@@ A.3 Training details
For training, we utilize 15,249 programming task instances provided in training split of TACO dataset, where we exclude instances without a ground-truth code solution and those whose unit test annotation do not follow stdio format (e.g., functional assertions) or without ground-truth code solution annotation, and 1,000 task instances for validation. We train Qwen3-4B and Qwen3-14B via UTRL. For Qwen3-4B experiment, we train the unit test generator for 50 steps in the first iteration, and train the code generator for 370 steps, and then continually train the unit test generator for 100 steps in the second iteration. For Qwen3-14B experiment, we train the unit test generator for 100 steps at the first iteration, without further iteration. We use weighted sum of discrimination reward and validity reward (see Equation 3) measured on the 1,000 validation tasks as a validation metric to choose training steps. For the number of LLM-generated code samples to define a discrimination reward, we set the value as 8. For $\tau$, a hyperparameter to adjust the desired minimum number of the test cases in the unit test, we set the value as 12. Additionally, for $\lambda$, a weighting factor for the discrimination reward, we set this value as 0.85.
@@ A.4 Evaluation details
Table 5: Distribution of evaluation tasks over multiple competitive programming platforms.
@@
For measuring Best-of-N improvement, we utilize LLM for sampling 32 code solutions, and select the best code solution by using the generated unit test as a code evaluator (i.e., select the code solution that passes the largest number of test cases in the generated unit test). We then evaluate code score and code accuracy regarding the selected code solution, using ground-truth unit tests provided in the TACO evaluation dataset. An intuition behind this evaluation is that a more discriminative unit test will identify and select higher-quality code among multiple candidate code solutions. We evaluate the Best-of-$N$ improvement on multiple code generator LLMs, Qwen3-4B, 8B, 14B, and GPT-4o.

### v3

@@ 5 Conclusion
Acknowledgement
This research was conducted as part of the Sovereign AI Foundation Model Project(Data Track), organized by the Ministry of Science and ICT(MSIT) and supported by the National Information Society Agency(NIA), S.Korea. (Grant No. 2026-AIData-WII01). This work was also supported by Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)). This work was also supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00414822).
@@ A.1 Grouped Relative Policy Optimization
A.2 Additional experiments
In this section, we provide supplementary experiments to thoroughly validate each component of UTRL, and demonstrate further advantage of UTRL.
A.2.1 Evaluation on LiveCodeBench
To evaluate whether the unit test generator LLM trained via UTRL generalizes to arbitrary diverse programming tasks, we additionally measure Best-of-N improvement and unit test fidelity of UTRL on 511 programming tasks in LiveCodeBench (Jain et al., 2024b). As shown in the Table 3, Qwen3-4B trained via UTRL achieves best-of-N improvement higher than GPT-4.1, and Qwen2.5-7B-Instruct trained via UTRL outperforms the same model trained via CURE. This results underscore that the effectiveness of unit test generator trained via UTRL well generalizes to diverse programming tasks.
Unit test generated by Code LLM: Qwen3-4B Code LLM: Qwen3-8B
Score Acc (%) Score Acc (%)
Qwen3-4B 0.704 55.2 0.714 55.4
Qwen3-4B + UTRL 0.744 59.9 0.752 59.3
Qwen2.5-7B 0.607 44.6 0.617 44.0
Qwen2.5-7B+CURE 0.660 48.2 0.671 47.7
Qwen2.5-7B+UTRL 0.686 54.4 0.693 52.4
GPT-4.1 0.732 58.3 0.735 58.3
Table 3: Best-of-N improvement evaluation on LiveCodeBench. A unit test generator trained via UTRL also generalizes to unseen tasks in LiveCodeBench, outperforming CURE. Moreover, Qwen3-4B trained via UTRL achieves performance superior to GPT-4.1.
A.2.2 Comparison to best-of-N distillation
As an additional baseline for training unit test generator, we compare UTRL with best-of-N distillation approach. For Best-of-N distillation, we sample 8 unit tests for each training instance using Qwen3-4B, select one with maximum discrimination reward, and finetune the Qwen3-4B with the selected unit test via SFT. As shown in the Table 4, although the best-of-N distillation approach yields performance comparable to SFT w/ $\mathcal{D}_{\text{UT}}$ baseline even without requiring ground-truth unit tests for training, UTRL achieves superior performance.
Unit test generated by Code LLM: Qwen3-4B Code LLM: Qwen3-8B
Score Acc (%) Score Acc (%)
Qwen3-4B + SFT w/ $\mathcal{D}_{\text{UT}}$ 0.457 11.7 0.458 11.7
Qwen3-4B + best-of-N distil. 0.423 10.2 0.437 9.9
Qwen3-4B + UTRL 0.555 14.5 0.578 14.9
Table 4: Comparison to best-of-N distllation approach for training unit test generator.
A.2.3 Experiment on different LLM
Although we mainly train Qwen3-4B and Qwen3-14B via UTRL, we additionally show effectiveness of UTRL in training arbitrary LLMs. To this end, we train Llama3.1-8B-Instruct via UTRL for 50 steps and evaluate best-of-N improvement of the unit tests generated by the resulting unit test generator. As shown in the Table 7, UTRL effectively enhances the unit test generation capability when applied to the Llama3.1-8B-Instruct, improving achieving best-of-N improvement of 11.75% on average, which is remarkably higher than Llama-3.1-8B-Instruct model.
Unit test generated by Code LLM: Qwen3-4B Code LLM: Qwen3-8B
Score Acc (%) Score Acc (%)
Llama-3.1-8B-Instruct 0.368 7.8 0.380 7.7
Llama-3.1-8B-Instruct + UTRL 0.475 11.6 0.494 11.9
Table 5: Performance of Llama-3.1-8B-Instruct trained via UTRL. UTRL effectively improve the unit test generation capability of Llama-3.1-8B-Instruct, which is measured by best-of-N improvement.
A.2.4 Ablation on weighting term for discrimination reward
We conduct ablation study on hyperparameter $\lambda$, which controls the weight of the discrimination reward for training unit test generator LLM. We sweep over $\lambda\in(0.0,0.85,1.0)$, and trained unit test generator for 50 steps for each options. As shown in the Table 6 and Table 7, setting $\lambda=0.0$ leads to improved validity (proportion of the functionally correct test cases among entire test cases in the unit test), while resulting in lower Best-of-N improvement, since the model is guided to generate test cases that are merely valid but indifferent to discriminativeness. Conversely, setting $\lambda=1.0$ results in a substantial fraction of invalid test cases, which also degrades unit test quality measured by Best-of-N improvement. Setting $\lambda=0.85$ achieves the best unit test quality, balancing validity and discriminativeness of unit tests. These results imply that both discrimination reward and validity reward are essential, and should be balanced appropriately to produce high-quality unit tests.
Code LLM: Qwen3-4B Code LLM: Qwen3-8B
Score Acc (%) Score Acc (%)
$\lambda=0$ 0.509 12.4 0.524 13.1
$\lambda=0.85$ 0.514 13.7 0.534 13.8
$\lambda=1.0$ 0.485 12.3 0.492 11.5
Table 6: Best-of-N improvement according to weighting factor for discrimination reward $\lambda$. Omitting validity reward ($\lambda=0$) degrades the quality of the unit tests owing to the functionally incorrect test cases occupying more than half of the generated unit test.
validity ratio (%)
$\lambda=0$ 64.3
$\lambda=0.85$ 57.2
$\lambda=1.0$ 49.7
Table 7: Ratio of valid test cases according to weighting factor for discrimination reward $\lambda$.
A.2.5 Ablation on validity reward
@@
In order to demonstrate the effectiveness of the validity reward and its design choice (see Equation 2), we conduct an ablation study comparing the following variants of the UTRL, (1) UTRL trained without validity reward (i.e., w/o $R_{\text{valid}}$), and (2) UTRL with validity reward but without clipped normalization in validity reward term, where the validity reward is defined as a ratio of valid test cases among the generated test cases (i.e., w/o clipping). Figure 10 shows the number of test cases and the ratio of valid test cases in the generated unit test over training steps. UTRL without $R_{\text{valid}}$ results in the generation of unit tests with more than 50% of invalid test cases, while UTRL results in unit tests containing less than 35% of invalid test cases. Furthermore, in the case of UTRL without clipping, we observe that the model collapses to generate an extremely small number of trivial test cases (i.e., replicating 2 3 basic test cases included in the problem statement) in order to maximize the ratio of the valid test cases, hindering the learning progress toward generating comprehensive unit tests. These results demonstrates the effectiveness of the design choice of UTRL.
A.2.6 Best-of-N improvement measured with additional code generators
@@
Qwen3-4B + UTRL 0.555 +0.207 14.5 +7.0 0.568 +0.353 16.9 +10.2
Qwen3-14B + UTRL 0.579 +0.219 15.8 +7.8 0.578 +0.359 17.0 +11.0
GT (upperbound) 0.630 +0.284 20.2 +13.0 0.599 +0.197 21.7 +11.0
Table 8: Best-of-N improvement measured on TACO evaluation set using Qwen3-4B and GPT-4o as code generator LLM. UTRL shows the highest best-of-N improvement compared to baselines. Here, UTRL denotes the variant of UTRL using 4.5K programming tasks for training, for fair comparison to CURE.
@@ A.3 Implementation details
For training, we utilize 15,249 programming task instances provided in training split of TACO dataset, where we exclude instances without a ground-truth code solution and those whose unit test annotation do not follow stdio format (e.g., functional assertions), and 1,000 programming task instances for validation. In our experiment, we train instruction-finetuned Qwen3-4B and Qwen3-14B via UTRL. For Qwen3-4B experiment, we train the unit test generator for 50 steps in the first iteration, and train the code generator for 330 steps, and then continually train the unit test generator for 170 steps in the second iteration. For Qwen3-14B experiment, we train the unit test generator for 100 steps at the first iteration, without further iteration. We use weighted sum of discrimination reward and validity reward (see Equation 3) measured on the 1,000 validation tasks as a validation metric to choose training steps. For the number of LLM-generated code samples to define a discrimination reward, we set the value as 8. For $\tau$, a hyperparameter to adjust the desired minimum number of the test cases in the unit test, we set the value as 12. Additionally, for $\lambda$, a weighting factor for the discrimination reward, we set this value as 0.85. Moreover, in order to ensure fair comparison with CURE, we also implement UTRL by training Qwen2.5-7B-Instruct via UTRL for 100 steps using the 4.5K instruction-code solution pairs in CodeContests dataset, following the experimental setup of CURE. For UTRL , we trained the model only for a single iteration.
@@ A.4 Evaluation details
Table 10: Distribution of TACO evaluation tasks over multiple competitive programming platforms.
@@
For measuring Best-of-N improvement, we utilize LLM for sampling 32 code solutions, and select the best code solution by using the generated unit test as a code evaluator (i.e., select the code solution that passes the largest number of test cases in the generated unit test). We then evaluate code score and code accuracy regarding the selected code solution, using ground-truth unit tests provided in the TACO evaluation dataset. An intuition behind this evaluation is that a more discriminative unit test will identify and select higher-quality code among multiple candidate code solutions. In TACO evaluation set, we evaluate the Best-of-$N$ improvement on multiple code generator LLMs, Qwen3-4B, 8B, 14B, and GPT-4o. Additionally, in LiveCodeBench-v2 evaluation set, we evaluate the Best-of-$N$ improvement on Qwen3-4B and Qwen3-8B code generator LLMs.